OpenStudy (anonymous):

Can anyone help me with this question! Create a rational expression to be your game piece. You may choose from the list of factors below or make your own. There must be a variable term in both the numerator and denominator. (5x) (2x) (x + 4) (x – 5) (2x + 1) (3x + 5)

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

Ok! I understand that but like i just need an equation thats all! @redbullx

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

yup and here it is! @redbullx

OpenStudy (anonymous):

thanks the picture needed for like the process Turn one. Flip your coin and perform the appropriate operation. Explain to the game master how to add your rational expression to the one on the correct space. Use complete sentences. @redbullx

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

Ok! @redbullx

OpenStudy (anonymous):

So you want me to solve it and then post my answer here afterwords? @redbullx

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@ranga this is basically the whole assignment. see the attachment below!

OpenStudy (ranga):

1. Painting a room can be done faster if two people share the work. 2. You can point a room in 4 hours if you work alone. Your friend, Mary, can paint the same room in 5 hours if she paints the room all by herself. How long will it take if you and Marie both paint the room together?

OpenStudy (ranga):

3. In one hour, you, Gaby, can paint 1/4 th of the room. In one hour, Mary can paint 1/5 th of the room. Assume it takes x hours to paint the room if Gaby and Marie both worked together. In one hour, Gaby and Marie can paint 1/x th of the room. Therefore, 1/4 + 1/5 = 1/x

OpenStudy (ranga):

4. 1/4 + 1/5 = 1/x 5/20 + 4/20 = 1/x 9/20 = 1/x x = 20/9 \[x = 2\frac{ 2 }{ 9 } hrs.\]

OpenStudy (anonymous):

OMG Thank you so much! You are a true help! @ranga

OpenStudy (ranga):

You are welcome.

OpenStudy (anonymous):

@austinL look at the attached file that i just posted!

OpenStudy (anonymous):

@austinL this is the second part!